Distance from a Point to a Line
8.Find the distance from the point P(3,5)to the line x=-2.
thank you!!!

Answer by user_dude2008(1862)

In this case,take difference from x-coord of P(3,5) and x=-2: 3-(-2)=3+2=5

Ans: distance from P(3,5) to line x=-2: 5 units


Graph to verify